The bookkeeper of TGIT Corporation had its May financial information on post-it-notes that were stuck to a computer monitor. The managers son was visiting the office. He was attracted to the brightly coloured pieces of paper and proceeded to colour on them with equally brightly coloured markers. Figures were obliterated on some of the post-it notes. The following information was retrieved Selling price per unit: $12 per unit Units sold: 20,000 units Variable cost per unit: $4.80 per unit Fixed costs: $54,000 Help reconstruct the May data by providing the value for each of the unknowns listed below: a) Total sales revenue: $ b) Total variable cost: $ c) Contribution margin per unit: $ d) Total contribution margin: S e) Operating income (income before taxes): $ f) Break even sales in units g) Variable cost ratio expressed in percent: h) Contribution margin ratio in percent: ) Break even sales in revenue dollars: $ k) The company wants a target profit before taxes of $125,000 in June? per unit units What is the total contribution margin that the company would require in June to hit its target? How many units would the company need to sell in June to hit its target in k)? units
Answer #1

a) Total sales revenue = 20000*12 = $240000

b) Total variable Cost = 20000*4.8 = $96000

c) Unit Contribution margin = 12-4.8 = $7.2 per unit

d) Total Contribution margin = 7.2*20000 = $144000

Note : Please post whole question in the group of 4-4 requirements as per HOMEWORKLIB RULES